Solve the substitution:
(3x + 5y = 5
(x -2y = 9

Answer by oscargut(686)

3x + 5y = 5
x -2y = 9 then x=2y+9
substituting in the first eq we have 3(2y+9)+5y=5
6y+27+5y=5
11y=-22
then y=-2 and x=2(-2)+9=5
Answer: x=5 and y=-2